# On the eigenvalue problem for a bulk/surface elliptic system

### v1

@@ 4. Proofs
Moreover, since $H$ is a rotation, the metric $(\cdot,\cdot)_{\Gamma}$ has the same matrix representation in the spherical coordinates $(y^{1},\ldots,y^{N_{1}})$ on $\Gamma_{1}=\partial B_{R_{1}}$, given by $\psi:\partial B_{R_{1}}\to\mathbb{R}^{N-1}$, and in the spherical coordinates on $\Gamma_{1}$ given by the map $\psi\cdot H:\partial B_{R_{1}}\to\mathbb{R}^{N-1}$. Hence, in these two local coordinates sytems, the volume element $d\mathcal{H}^{N-1}=\sqrt{g}dy^{1}\wedge\ldots\wedge dy^{N-1}$, the Riemannian gradient $\nabla_{\Gamma}$ and $|\cdot|_{\Gamma}|^{2}$ are equals. Hence, since $H:\Gamma_{1}\to\Gamma_{1}$ is bijective and $C^{\infty}$, we have
@@
Now let $u$ be a real–valued weak solution of (16). We claim that it is radially symmetric. If $u$ is trivial there is nothing to prove. Hence, using Theorem 4, we can suppose that $u$ is positive in $\Omega\cup\Gamma_{1}$. By (61), (64), the minimality asserted in Theorem 2 and the simplicity of the $\lambda_{1}$ asserted in Theorem 4, we get that for any rotation $H$ there is $\mu=\mu_{H}\in\mathbb{R}$ such that $u\cdot H=\mu_{H}u$ in $\Omega$. By (64) we also have $\mu_{H}\in\{-1,1\}$. Since $u$ is positive then $\mu_{H}=1$, so $u\cdot H=u$ in $\Omega$. Since $H$ is arbitrary our claim is proved. Passing to real and imaginary parts all weak solutions of (16) are radially symmetric.
@@
(65) $\begin{cases}-w^{\prime\prime}(\rho)-\frac{N-1}{\rho}w^{\prime}(\rho)=\lambda_{1}w(\rho)\quad&\text{in $(R_{1},R_{2})$,}\\ -w^{\prime}(R_{1})=\lambda_{1}w(R_{1}),\quad&u(R_{2})=0.\end{cases}$

### v2

@@ 4. Proofs
Moreover, since $H$ is a rotation, the metric $(\cdot,\cdot)_{\Gamma}$ has the same matrix representation in the spherical coordinates $(y^{1},\ldots,y^{N-1})$ on $\Gamma_{1}=\partial B_{R_{1}}$, given by $\psi:\partial B_{R_{1}}\to\mathbb{R}^{N-1}$, and in the spherical coordinates on $\Gamma_{1}$ given by the map $\psi\cdot H:\partial B_{R_{1}}\to\mathbb{R}^{N-1}$. Hence, in these two local coordinates sytems, the volume element $d\mathcal{H}^{N-1}=\sqrt{g}dy^{1}\wedge\ldots\wedge dy^{N-1}$, the Riemannian gradient $\nabla_{\Gamma}$ and $|\cdot|_{\Gamma}|^{2}$ are equals. Hence, since $H:\Gamma_{1}\to\Gamma_{1}$ is bijective and $C^{\infty}$, we have
@@
Now let $u$ be a real–valued weak solution of (16). We claim that $u$ is radially symmetric. If $u$ is trivial there is nothing to prove. Hence, using Theorem 4, we can suppose that $u$ is positive in $\Omega\cup\Gamma_{1}$. By (61), (64), the minimality asserted in Theorem 2 and the simplicity of the $\lambda_{1}$ asserted in Theorem 4, we get that for any rotation $H$ there is $\mu=\mu_{H}\in\mathbb{R}$ such that $u\cdot H=\mu_{H}u$ in $\Omega$. By (64) we also have $\mu_{H}\in\{-1,1\}$. Since $u$ is positive then $\mu_{H}=1$, so $u\cdot H=u$ in $\Omega$. Since $H$ is arbitrary our claim is proved. Passing to real and imaginary parts all weak solutions of (16) are radially symmetric.
@@
(65) $\begin{cases}-w^{\prime\prime}(\rho)-\frac{N-1}{\rho}w^{\prime}(\rho)=\lambda_{1}w(\rho)\quad&\text{in $(R_{1},R_{2})$,}\\ -w^{\prime}(R_{1})=\lambda_{1}w(R_{1}),\quad w(R_{2})=0.&\end{cases}$